Use Cramer’s rule to solve for x, y and z

2x + y − 3z = 0

4x + 5y + z = 4

x + y − 4z = −1 


1
Expert's answer
2021-06-01T12:59:06-0400

First we find the coefficient determinant:

D="\begin{vmatrix}
 2 & 1&-3 \\
 4 & 5&1\\
1&1&-4
\end{vmatrix}" =-40-12+1+15-2+16=-22 .

Then we form and evaluate Dx by replacing the first column of values with the answer column:

Dx = "\begin{vmatrix}
 0& 1&-3 \\
 4 & 5&1\\
-1&1&-4
\end{vmatrix}" = 0-12-1-15+0+16=-12.

Form and evaluate Dy by replacing the second column of values with the answer column:

Dy = "\begin{vmatrix}
 2& 0&-3 \\
 4 & 4&1\\
1&-1&-4
\end{vmatrix}" = -32+12+0+12+2+0=-6 .

Form and evaluate Dz by replacing the third column of values with the answer column:

Dz = "\begin{vmatrix}
 2& 1&0 \\
 4 & 5&4\\
1&1&-1
\end{vmatrix}" = -10+0+4-8+0+4=-10.

Cramer's rule says that x=Dx/D; y=Dy/D; z=Dz/D.

That is:

x=-12/-22=6/11; y=-6/-22=3/11; z=-10/-22=5/11.

The final answer:

(x,y,z)=(6/11, 3/11, 5/11) .